# UNEM: UNrolled Generalized EM for Transductive Few-Shot Learning

### v1

@@ Abstract
Transductive few-shot learning has recently triggered wide attention in computer vision. Yet, current methods introduce key hyper-parameters, which control the prediction statistics of the test batches, such as the level of class balance, affecting performances significantly. Such hyper-parameters are empirically grid-searched over validation data, and their configurations may vary substantially with the target dataset and pre-training model, making such empirical searches both sub-optimal and computationally intractable. In this work, we advocate and introduce the unrolling paradigm, also referred to as “learning to optimize”, in the context of few-shot learning, thereby learning efficiently and effectively a set of optimized hyper-parameters. Specifically, we unroll a generalization of the ubiquitous Expectation-Maximization (EM) optimizer into a neural network architecture, mapping each of its iterates to a layer and learning a set of key hyper-parameters over validation data. Our unrolling approach covers various statistical feature distributions and pre-training paradigms, including recent foundational vision-language models and standard vision-only classifiers. We report comprehensive experiments, which cover a breadth of fine-grained downstream image classification tasks, showing significant gains brought by the proposed unrolled EM algorithm over iterative variants. The achieved improvements reach up to 10% and 7.5% on vision-only and vision-language benchmarks, respectively. The source code and learned parameters are available at https://github.com/ZhouLong0/UNEM-Transductive.
@@ 1 Introduction
Several recent attempts tackled this limitation, to handle more realistic scenarios [48, 32, 46, 24, 33]. Yet, these methods introduce key hyper-parameters, which control the prediction statistics of the unlabeled query set, such as the level of class balance. Such hyper-parameters are empirically grid-searched over pre-defined sets of values using validation data, and their optimal configurations may vary substantially with the target dataset and pre-training model [17]. To illustrate this, we depict in Fig. 1 the accuracy as a function of a class-balance hyper-parameter for the recent method in [33]. One may observe that this hyper-parameter has a crucial effect on the performance, and its optimal value might vary by orders of magnitude across the target datasets. The issue is further compounded by additional hyper-parameters, which makes intensive empirical grid searches for the hyper-parameters over validation data and pre-defined intervals of values both sub-optimal and computationally intractable. Therefore, we advocate and introduce the unrolling paradigm (also called “learning to optimize”) in the context of few-shot learning, which enables to learn efficiently and effectively a set of optimized hyper-parameters. Specifically, our main contributions could be summarized as follows:
@@
    We unroll the generalized EM optimizer into a neural network architecture, mapping each of its iterates to a layer and learning the introduced hyper-parameters over validation data. Our unrolling approach offers greater flexibility in optimizing the hyper-parameters, allowing them to vary across the network’s layers. To the best of our knowledge, this study is the first to investigate unrolling in transductive few-shot learning.
@@
    We report comprehensive experiments, which cover a breadth of fine-grained downstream image classification tasks and different pre-training models, showing consistent and substantial gains over recent state-of-the-art methods.
@@ 2 Related works
Few-shot classification with vision-language models: Contrastive Vision-Language Pre-training (CLIP) has recently emerged as an effective model for enhancing various vision tasks through visual-text pairs. By learning transferable visual models with natural language supervision, CLIP demonstrates strong performance in zero-shot classification by matching the image features to text embeddings of novel classes [38]. To further enhance its classification capabilities, several studies have extended CLIP to the few-shot setting [18, 55, 6, 53, 56]. While the linear probe model [38] trains a logistic regression classifier using CLIP image features, the authors of proposed in [18] proposed a generalization of this baseline, modeling the classifier weights as learnable functions of the visual prototypes and text embeddings. Context Optimization (CoOp) [55] and its extended versions [53, 6, 56] have been developed based on the concept of prompt learning. While CoOp [55] aims to model context in prompts using continuous representations, a knowledge-guided CoOp is proposed in [53] to enhance the generalization ability by minimizing the discrepancy between the learnable prompts and the original ones. In [6], the authors learn multiple prompts, describing the characteristics of each class, through the minimization of an optimal-transport distance. Unlike prompt learning methods, which fine-tune the input text, another family of methods, referred to as adapters, aims to transform the visual or text encoders, such as [54, 12]. For instance, TIP-Adapter [54] adds a non-parametric adapter to the weight-frozen CLIP model, and updates the prior knowledge encoded in CLIP by feature retrieval. CLIP-Adapter [12] introduces a multi-layered perceptron to learn new features, and combine them with the original CLIP-encoded features via residual connections.
@@
Class-balance and hyperparameters setting: Most of the existing transductive few-shot methods have been designed for perfectly balanced query sets (i.e., uniform class distribution), and have shown drops in performances under class-imbalanced settings. To address this flaw, various methods have been developed in the last years [48, 32, 46, 24, 33]. For instance, In [46], the categorical probability of each query sample is regularized to quantify the difference between the class marginal distribution and the uniform one. The works in [48, 32, 33, 24] explored various weighted terms, which are added to the objective functions, to mitigate the effect of the class-balance bias. To this end, some hyper-parameters are introduced, to weigh to contributions of such terms in the objective function. Unlike [32], where the introduced hyper-parameter has been theoretically chosen in order to compensate for the hidden class-balance bias in the used clustering objective, the hyper-parameters are often set in an empirical manner based on the validation classes of each dataset. Thus, a set of hyperparameter values (in a given range) are evaluated, and the ones yielding the highest accuracy are selected.
@@ 3.1 Generalized EM algorithm
  • The second term in (1) controls the partition complexity of the model, penalizing the number of non-empty clusters in the solution. It corresponds to the Shannon entropy of class distribution $(\pi_{k})_{1\leq k\leq K}$, defined as:
@@
  • The third term in (1) is an entropic barrier, enabling to soften assignments $u_{n,k}$, while imposing a non-negativity constraint on each of them. It is given by:
@@
     // Update Distribution parameters for each class using a given an estimation algorithm (denoted here by “DP_est”)
@@
  • The first one is the Gaussian distribution, which is commonly used in standard clustering and transductive few-shot-methods applied to vision-only models [44, 48, 32]. By assuming a Gaussian distribution with mean $\boldsymbol{\theta}_{k}$ and identity covariance matrix, the pdf $\mathrm{p}(\boldsymbol{z}_{n}|\boldsymbol{\theta}_{k})$ reads as follows:
@@
  • The second one is the Dirichlet distribution, which has recently shown good modelling performance in the context of transductive few-shot for vision-language models such as CLIP [33]. For $\boldsymbol{z}_{n}=(z_{n,i})_{1\leq i\leq K}$ and $\boldsymbol{\theta}_{k}=(\theta_{k,i})_{1\leq i\leq K}$, the associated pdf is given by:
@@
    where the normalization factor $\mathcal{B}(\boldsymbol{\theta}_{k})$ is:
@@ 3.2 UNrolled EM architecture (UNEM)
where $(\lambda^{(\ell)},T^{(\ell)})_{0\leq\ell\leq L-1}$ is the vector of hyper-parameters to be learned. This leads to the UNrolled EM (UNEM) model shown in Fig. 2. The latter shows the inputs and outputs of a given layer $\mathcal{L}^{(l)}$ as well as its three main blocks for the update rules.
@@
More specifically, to ensure the non-negativity of hyper-parameter $\lambda^{(\ell)}$, we propose to express it as a Softplus function, which could be seen as a smooth approximation of the RELU activation function, yielding:
@@ 4.1 Experimental settings
Task generation: We adopt a realistic transductive few-shot evaluation protocol, which is in line with state-of-the-art evaluation protocols [33, 32, 43, 48, 29]. While $K$ designates the total number of classes in the labeled support set $\mathbb{S}$, $K_{\rm eff}$ (with $K_{\rm eff}\ll K$) denotes the number of effective classes present in the unlabeled query set $\mathbb{Q}$. The classes in the query set remain undisclosed during inference, while randomly selecting $|\mathbb{Q}|$ samples. On the other hand, the support set is built by uniformly selecting $s$ images from each of the $K$ classes. In this paper, the few-shot tasks are performed with (i) 5, 10, and 20 shots for vision-only models, and (ii) 4 shots for vision-language models. Moreover, in both of our evaluation scenarios, we used $K_{\rm eff}=5$ and $|\mathbb{Q}|=75$.
@@
On the other hand, the training of our unrolled architecture, composed of 10 layers (i.e., $L=10$), is performed on several tasks sampled from the validation set of each dataset. This implies that the proposed solution is highly economical in terms of parameters, typically requiring only a few tenths, compared to standard neural network architectures. Let us recall that the validation set is often used to select the hyperparameters in recent state-of-the-art transductive methods as mentioned in Section 2. Specifically, for the vision-only (resp. vision-language) models, the training of the unrolled architecture is carried out on 1,000 (resp. 100) tasks, while using an initial learning rate of 0.1 (resp. 0.5) with a decay factor of 0.5, a number of epochs equals to 80, and ADAM as optimizer. Our architecture is implemented in Pytorch (version 2.3.0) and run on NVIDIA QUADRO RTX8000 (with 48 GB of memory).
@@
Ind. Tip-Adapter [54] 76.7 72.5 54.7 86.4 83.2 88.8 72.1 23.7 63.9 66.7 68.9
CoOp [55] 76.3 63.2 52.2 86.2 81.0 87.7 67.0 22.2 61.3 63.4 66.1
Transd. BDSCPN [28] 74.7 46.1 45.2 81.3 74.2 82.0 59.0 18.0 48.1 54.5 58.3
Laplacian Shot [58] 76.6 53.0 52.6 88.4 85.5 86.8 67.0 22.2 60.4 63.8 65.6
$\alpha$-TIM [48] 66.1 46.1 45.3 87.1 79.1 83.3 59.4 20.4 53.4 53.4 59.4
PADDLE [32] 71.8 45.9 50.0 84.7 82.3 81.9 63.7 21.3 56.1 60.6 61.6
EM-Dirichlet [33] 88.7 50.8 62.6 92.5 91.3 90.1 76.1 24.9 73.5 80.9 73.1
UNEM-Dirichlet 91.4 53.8 65.3 96.0 95.6 93.4 78.5 30.4 80.0 88.5 77.3
Table 3: Comparison of the proposed UNEM-Dirichlet with respect to the state-of-the-art methods on 10 different datasets. The metric is accuracy (in percentage). Results are averaged across 1,000 tasks.
@@ 4.3 UNEM-Dirichlet for few-shot CLIP
Datasets: The second unrolled architecture, designated by UNEM-Dirichlet, is assessed on 10 benchmark datasets that are commonly used for CLIP scenario: Caltech101 [10], OxfordPets [36], StanfordCars [22], Flowers102 [35], Food101 [3], FGVCAircraft [31], SUN397 [52], DTD [9], EuroSAT [15], and UCF101 [45]. These datasets cover diverse classification challenges, from object recognition (Caltech101, Food101) to fine-grained tasks (StanfordCars, FGVCAircraft) and scene understanding (SUN397, EuroSAT). For each dataset, the vision-text feature vectors $\boldsymbol{z}_{n}$ are extracted using the CLIP’s pre-trained model, while applying a temperature parameter $T_{z}$ as described in Appendix C. This temperature parameter has also been learned through the unrolling approach.
@@
Appendix D includes additional results to illustrate, in both scenarios, the effects of the temperature parameter in this framework, and to show the benefits of learning variable hyper-parameters across the architecture layers.
@@ 5 Conclusion
In this paper, we focus on transductive few-shot methods grounded on a generalized form of the EM algorithm. These methods include a parameter to control class balance. The proposed GEM algorithm, applicable to any mixture of distributions, incorporates also a temperature scaling parameter. The optimization process is then unrolled into a neural network architecture, enabling efficient learning of the introduced hyper-parameters. The results demonstrate the effectiveness of the proposed approach on both vision-only and vision-language models. In future work, it would be valuable to explore the potential of unrolling techniques across a broader range of computer vision tasks, especially with the recent rise of foundational vision-langauge models.
@@ D.1 Ablation studies
In this part, we perform ablation studies to illustrate the effects of the introduced temperature scaling parameter and show the benefits of learning adaptive hyper-parameters across the unrolled architecture layers.
@@
Tables 4 and 5 depict the accuracy results in vision-only few-shot setting. Thus, it can be noticed that including the temperature scaling yields an accuracy improvement, which may vary from 1% to 3%. Moreover, in the context of vision-language models whose accuracy results are shown in Table 6, similar gains (reaching up to 3%), depending on the target donwstream dataset, are also achieved. This confirms again the advantage of incorporating the temperature scaling in our generalized algorithm.
@@
Tables 7 and 8 provide the accuracy results for fixed and adaptive hyper-parameters optimization with vision-only models. It can be seen that learning adaptive hyper-parameters yields an accuracy gain of about 2-4% compared to the case when the hyper-parameters are kept fixed across layers. Similar comparisons are also performed with vision-language models as shown in Table 9. In this context, the improvement achieved by learning adaptive hyper-parameters often ranges from 1 to 2%.
@@
D.2 Illustration and analysis of the learned hyperparameters

### v2

@@ Abstract
Transductive few-shot learning has recently triggered wide attention in computer vision. Yet, current methods introduce key hyper-parameters, which control the prediction statistics of the test batches, such as the level of class balance, affecting performances significantly. Such hyper-parameters are empirically grid-searched over validation data, and their configurations may vary substantially with the target dataset and pre-training model, making such empirical searches both sub-optimal and computationally intractable. In this work, we advocate and introduce the unrolling paradigm, also referred to as “learning to optimize", in the context of few-shot learning, thereby learning efficiently and effectively a set of optimized hyper-parameters. Specifically, we unroll a generalization of the ubiquitous Expectation-Maximization (EM) optimizer into a neural network architecture, mapping each of its iterates to a layer and learning a set of key hyper-parameters over validation data. Our unrolling approach covers various statistical feature distributions and pre-training paradigms, including recent foundational vision-language models and standard vision-only classifiers. We report comprehensive experiments, which cover a breadth of fine-grained downstream image classification tasks, showing significant gains brought by the proposed unrolled EM algorithm over iterative variants. The achieved improvements reach up to 10% and 7.5% on vision-only and vision-language benchmarks, respectively. The source code and learned parameters are available at https://github.com/ZhouLong0/UNEM-Transductive.
@@ 1 Introduction
Several recent attempts tackled this limitation, to handle more realistic scenarios [50, 32, 48, 24, 33]. Yet, these methods introduce key hyper-parameters, which control the prediction statistics of the unlabeled query set, such as the level of class balance. Such hyper-parameters are empirically grid-searched over pre-defined sets of values using validation data, and their optimal configurations may vary substantially with the target dataset and pre-training model [17]. To illustrate this, we depict in Fig. 1 the accuracy as a function of a class-balance hyper-parameter for the recent method in [33]. One may observe that this hyper-parameter has a crucial effect on the performance, and its optimal value might vary by orders of magnitude across the target datasets. The issue is further compounded by additional hyper-parameters, which makes intensive empirical grid searches for the hyper-parameters over validation data and pre-defined intervals of values both sub-optimal and computationally intractable. Therefore, we advocate and introduce the unrolling paradigm (also called “learning to optimize") in the context of few-shot learning, which enables to learn efficiently and effectively a set of optimized hyper-parameters. Specifically, our main contributions could be summarized as follows:
@@
    We unroll the generalized EM optimizer into a neural network architecture, mapping each of its iterations to a layer and learning the introduced hyper-parameters over validation data. Our unrolling approach offers greater flexibility in optimizing the hyper-parameters, allowing them to vary across the network’s layers. To the best of our knowledge, this study is the first to investigate unrolling in transductive few-shot learning.
@@
    We report comprehensive experiments, which cover a breadth of fine-grained downstream image classification tasks and different pre-training models, showing substantial gains over recent state-of-the-art methods.
@@ 2 Related works
Few-shot classification with vision-language models: Contrastive Vision-Language Pre-training (CLIP) has recently emerged as an effective model for enhancing various vision tasks through visual-text pairs. By learning transferable visual models with natural language supervision, CLIP demonstrates strong performance in zero-shot classification by matching the image features to text embeddings of novel classes [39]. To further enhance its classification capabilities, several studies have extended CLIP to the few-shot setting [18, 57, 6, 55, 58]. While the linear probe model [39] trains a logistic regression classifier using CLIP image features, the authors of [18] proposed a generalization of this baseline, modeling the classifier weights as learnable functions of the visual prototypes and text embeddings. Context Optimization (CoOp) [57] and its extended versions [55, 6, 58] have been developed based on the concept of prompt learning. While CoOp [57] aims to model context in prompts using continuous representations, a knowledge-guided CoOp is proposed in [55] to enhance the generalization ability by minimizing the discrepancy between the learnable prompts and the original ones. In [6], the authors learn multiple prompts, describing the characteristics of each class, through the minimization of an optimal-transport distance. Unlike prompt learning methods, which fine-tune the input text, another family of methods, referred to as adapters, aims to transform the visual or text encoders, such as [56, 12]. For instance, TIP-Adapter [56] adds a non-parametric adapter to the weight-frozen CLIP model, and updates the prior knowledge encoded in CLIP by feature retrieval. CLIP-Adapter [12] introduces a multi-layered perceptron to learn new features, and combine them with the original CLIP-encoded features via residual connections.
@@
Class-balance and hyperparameters setting: Hyperparameter optimization has a large impact on the performance of a given iterative algorithm or learning model, and is generally tackled through greedy or gradient-based methods [34]. In the context of transductive few-shot learning, most of the existing methods have been designed for perfectly balanced query sets (i.e., uniform class distribution), and have shown drops in performances under class-imbalanced settings. To address this flaw, various methods have been developed in the last years [50, 32, 48, 24, 33]. For instance, in [48] the categorical probability of each query sample is regularized to quantify the difference between the class marginal distribution and the uniform one. The works in [50, 32, 33, 24] explored various weighted terms, which are added to the objective functions, to mitigate the effect of the class-balance bias. To this end, some hyper-parameters are introduced, to weigh to contributions of such terms in the objective function. Unlike [32], where the introduced hyper-parameter has been theoretically chosen in order to compensate for the hidden class-balance bias in the used clustering objective, the hyper-parameters are often set in an empirical manner based on the validation classes of each dataset. Thus, a set of hyperparameter values (in a given range) are evaluated, and the ones yielding the highest accuracy are selected.
@@ 3.1 Generalized EM algorithm
  • The second term in (1) controls the partition complexity of the model, penalizing the number of non-empty clusters in the solution. It corresponds to the Shannon entropy of class distribution $(\pi_{k})_{1\leq k\leq K}$, defined as
@@
  • The third term in (1) is an entropic barrier, enabling to soften assignments $u_{n,k}$, while imposing a non-negativity constraint on each of them. It is given by
@@
     // Update Distribution parameters for each class using a given estimation algorithm (denoted here by “DP_est”)
@@
  • The first one is the Gaussian distribution, which is commonly used in standard clustering and transductive few-shot-methods applied to vision models [46, 50, 32]. By assuming a Gaussian distribution with mean $\boldsymbol{\theta}_{k}$ and identity covariance matrix, the pdf $\mathrm{p}(\boldsymbol{z}_{n}|\boldsymbol{\theta}_{k})$ reads as follows
@@
  • The second one is the Dirichlet distribution, which has recently shown good modelling performance in the context of transductive few-shot for vision-language models such as CLIP [33]. For $\boldsymbol{z}_{n}=(z_{n,i})_{1\leq i\leq K}$ and $\boldsymbol{\theta}_{k}=(\theta_{k,i})_{1\leq i\leq K}$, the associated pdf is given by
@@
    where the normalization factor $\mathcal{B}(\boldsymbol{\theta}_{k})$ is
@@ 3.2 UNrolled EM architecture (UNEM)
where $(\lambda^{(\ell)},T^{(\ell)})_{0\leq\ell\leq L-1}$ is the vector of hyper-parameters to be learned. This leads to the UNrolled EM (UNEM) model shown in Fig. 2, which depicts the inputs and outputs of a given layer $\mathcal{L}^{(l)}$ as well as its three main blocks for the update rules.
@@
More specifically, to ensure the non-negativity of hyper-parameter $\lambda^{(\ell)}$, we propose to express it as a Softplus function, which could be seen as a smooth approximation of the RELU activation function, yielding
@@ 4.1 Experimental settings
Task generation: We adopt a realistic transductive few-shot evaluation protocol, which is in line with state-of-the-art protocols [33, 32, 45, 50, 29]. While $K$ designates the total number of classes in the labeled support set $\mathbb{S}$, $K_{\rm eff}$ (with $K_{\rm eff}\ll K$) denotes the number of effective classes present in the unlabeled query set $\mathbb{Q}$. The classes in the query set remain undisclosed during inference, while randomly selecting $|\mathbb{Q}|$ samples. On the other hand, the support set is built by uniformly selecting $s$ images from each of the $K$ classes. In this paper, the few-shot tasks are performed with (i) 5, 10, and 20 shots for vision-only models, and (ii) 4 shots for vision-language models. Moreover, in both of our evaluation scenarios, we used $K_{\rm eff}=5$ and $|\mathbb{Q}|=75$.
@@
On the other hand, the training of our unrolled architecture, composed of 10 layers (i.e., $L=10$), is performed on several tasks sampled from the validation set of each dataset. This implies that the proposed solution is highly economical in terms of parameters, typically requiring only a few tenths, compared to standard neural network architectures. Let us recall that the validation set is often used to select the hyperparameters in recent state-of-the-art transductive methods as mentioned in Section 2. Specifically, for the vision-only (resp. vision-language) models, the training of the unrolled architecture is carried out on 1,000 (resp. 100) tasks, while using an initial learning rate of 0.1 (resp. 0.5) with a decay factor of 0.5, a number of epochs equals to 80, and ADAM as optimizer. The trainable hyper-parameters are initialized to the default values set in [32] (i.e., $\lambda=|Q|$ for the vision-only models) and [33] (i.e., $\lambda=\frac{K}{K_{\rm eff}}$ for the vision-language models, with $T=1$). Our architecture is implemented in Pytorch (version 2.3.0) and run on NVIDIA QUADRO RTX8000 (with 48 GB of memory).
@@
Ind. Tip-Adapter [56] 76.7 72.5 54.7 86.4 83.2 88.8 72.1 23.7 63.9 66.7 62.7 68.3
CoOp [57] 76.3 63.2 52.2 86.2 81.0 87.7 67.0 22.2 61.3 63.4 59.9 65.5
Transd. BDSCPN [28] 74.7 46.1 45.2 81.3 74.2 82.0 59.0 18.0 48.1 54.5 49.2 57.5
Laplacian Shot [60] 76.6 53.0 52.6 88.4 85.5 86.8 67.0 22.2 60.4 63.8 56.3 64.8
$\alpha$-TIM [50] 66.1 46.1 45.3 87.1 79.1 83.3 59.4 20.4 53.4 53.4 42.7 57.8
PADDLE [32] 71.8 45.9 50.0 84.7 82.3 81.9 63.7 21.3 56.1 60.6 52.1 60.9
EM-Dirichlet [33] 88.7 50.8 62.6 92.5 91.3 90.1 76.1 24.9 73.5 80.9 78.4 73.6
UNEM-Dirichlet 91.4 53.8 65.3 96.0 95.6 93.4 78.5 30.4 80.0 88.5 83.1 77.8
Table 3: Comparison of the proposed UNEM-Dirichlet with respect to the state-of-the-art methods on 11 different datasets. The metric is accuracy (in percentage). Results are averaged across 1,000 tasks.
@@ 4.3 UNEM-Dirichlet for few-shot CLIP
Datasets: The second unrolled architecture, designated by UNEM-Dirichlet, is assessed on 11 benchmark datasets that are commonly used for CLIP scenario: Caltech101 [10], OxfordPets [37], StanfordCars [22], Flowers102 [36], Food101 [3], FGVCAircraft [31], SUN397 [54], DTD [9], EuroSAT [15], UCF101 [47], and ImageNet [42]. These datasets cover diverse classification challenges, from object recognition (Caltech101, Food101) to fine-grained tasks (StanfordCars, FGVCAircraft) and scene understanding (SUN397, EuroSAT). For each dataset, the vision-text feature vectors $\boldsymbol{z}_{n}$ are extracted using the CLIP’s pre-trained model, while applying a temperature parameter $T_{z}$ as described in Appendix C. This temperature parameter has also been learned through the unrolling approach.
@@
Appendix D includes additional results to illustrate, in both scenarios, the effects of the temperature parameter in this framework, and to show the benefits of learning variable hyperparameters across the architecture layers.
@@ 5 Conclusion
In this paper, we focus on transductive few-shot methods grounded on a generalized form of the EM algorithm. These methods include a parameter to control class balance. The proposed GEM algorithm, applicable to any mixture of distributions, incorporates also a temperature scaling parameter. The optimization process is then unrolled into a neural network architecture, enabling efficient learning of the introduced hyper-parameters. The results demonstrate the effectiveness of the proposed approach on both vision-only and vision-language models. In future work, it would be valuable to explore the potential of unrolling techniques across a broader range of computer vision tasks, especially with the recent rise of foundational vision-language models.
@@ D.1 Ablation studies
In this part, we perform ablation studies to illustrate the impact of the network depth, the effects of the introduced temperature scaling parameter and the benefits of learning adaptive hyper-parameters across the unrolled architecture layers.
$\bullet$ Impact of the unrolled architecture depth
First, we propose to analyze the impact of the number $L$ of layers of our unrolled architecture on the model accuracy, model size, and computational time. Table 4 reports the results. Thus, one of the main advantages of our UNEM model is that a few layers (about 7 or 10) are enough to achieve good performance. In what follows, the experiments are conducted using $L=10$.
#Layers ($L$) #Params Acc. Train Time (s) Inference Time/task (s)
3 7 65.6 2.80 $23.04\times 10^{-}$ 2
5 11 65.9 3.22 $23.09\times 10^{-}$ 2
7 15 66.3 3.46 $23.21\times 10^{-}$ 2
10 21 66.4 3.61 $23.36\times 10^{-}$ 2
12 25 66.2 4.06 $23.45\times 10^{-}$ 2
15 31 66.1 4.29 $23.48\times 10^{-}$ 2
18 37 66.2 5.74 $23.68\times 10^{-}$ 2
Table 4: Impact of the number of layers ($L$) on UNEM-Gaussian performance using mini-ImageNet, 5-shot and ResNet18.
@@
Tables 5 and 6 depict the accuracy results in vision-only few-shot setting. Thus, it can be noticed that including the temperature scaling yields an accuracy improvement, which may vary from 1% to 3%. Moreover, in the context of vision-language models whose accuracy results are shown in Table 7, similar gains (reaching up to 3%), depending on the target donwstream dataset, are also achieved. This confirms again the advantage of incorporating the temperature scaling in our generalized algorithm.
@@
Tables 8 and 9 provide the accuracy results for fixed and adaptive hyper-parameters optimization with vision-only models. It can be seen that learning adaptive hyper-parameters yields an accuracy gain of about 2-4% compared to the case when the hyper-parameters are kept fixed across layers. Similar comparisons are also performed with vision-language models as shown in Table 10. In this context, the improvement achieved by learning adaptive hyper-parameters often ranges from 1 to 2%.
@@
D.2 Performance under distribution shifts
In Table 11, we include the accuracy of the original PADDLE algorithm [32] and its unrolled version, with tiered-ImageNet used for pre-training, and a fine-grained classification dataset (CUB) as well as mini-ImageNet used for inference. One could observe similar improvements (about 4-7%) brought by UNEM.
Method CUB mini-ImageNet
PADDLE [32] 66.0 82.9
UNEM-Gaussian 72.9 87.0
Table 11: Cross-domain evaluation: Accuracy performance on mini-ImageNet and CUB using a model trained on tiered-ImageNet, with a 5-shot setting and a ResNet18 backbone.
D.3 Backbone effect in CLIP
The performance of EM-Dirichlet and UNEM-Dirichlet using ViT-B/32 as backbone are shown in Table 12. Thus, in comparison to CLIP with ResNet50 (see Table 3), one may observe a similar or slightly better accuracy performance with most datasets. Moreover, the gains brought by UNEM over the iterative variant are consistent with those observed with ResNet50 backbone.
Method
Food101
EuroSAT
DTD
OxfordPets
Flowers102
Caltech101
UCF101
FGVC Aircraft
Stanford Cars
SUN397
EM-Dirichlet [33] 89.3 54.8 63.9 92.5 92.7 92.8 77.2 27.3 73.9 81.7
UNEM-Dirichlet 91.4 57.5 67.4 95.7 95.2 94.0 80.4 33.8 77.8 88.4
Table 12: Accuracy performance of iterative EM-Dirichlet [33] and our UNEM variant using CLIP ViT-B/32 for feature extraction.
E Illustration and analysis of the learned hyperparameters